form Solve in matrix 1.) 3x + 6y eld Tx ty s
Answer #1

That's easy.

Thumb up.

Have a good day!!!

sol :- 3x+6y=10 - 2+= 5 . Using mateix form Ax=B 13 ][]-[5] A[34] *- [s], 8-[8] - [A: B] [ bo] C R 3 R₂ + R, . 1361107 Log :